OpenStudy (anonymous):

A VENDER PURCHASED 50 DOZEN BANANAS FOR RS600. 5 DOZEN BANANAS COLD NOT BE SOLD BECASE THEY GOT ROTTEN .AT WHAT PRICE PER DOZEN SHOLD THE VENDR SELL THE REMAINING BANANAS SO THAT HE MAKE AN OVER ALL PROFIT OF 20%

OpenStudy (shubhamsrg):

Profit = 20% i.e. he has to sell those 45 dozens for total of 600+ (20% of 600) = 720

OpenStudy (anonymous):

cost price of 50 doz bananas = 600 rs profit = 20% Total sale price = 120% of 600 = 600 * (120/100) = 720 rs Quantity available for sale = 50 - 5 = 45 doz sale price of banana = Total sale price / quantity = (720/45) = 16 rs per dozen So he should sell remaining 45 dozen good bananas at a price of Rs 16 / dozen to make a profit of 20%

OpenStudy (anonymous):

is this right

OpenStudy (shubhamsrg):

yep, seems right.